The production cost, represented by y, of smart phone devices are equal to the product of computing components a and the square of the number of devices x, added to the product pf labour b and number of devices x; after which savings on reusable material c are subtracted.

2.1. Suppose the production cost is R 239250, computing components cost is R1035, labour cost is R4140 and the savings on reusable materials is R1725, determine the number of devices produced

2.2. Hence, write down the equation with y as subject (in terms of x) to represent the production cost.( the value of a, b and c must be indicated.




1
Expert's answer
2022-03-01T10:59:45-0500

2.1) 

"y=ax^2+bx-c"


2.2


"y=239250"

"a=1035"

"b=4140"

"c=1725"

"1035x^2+4140x-240975=0, x>0"



"23x^2+92x-5355=0"


"x=\dfrac{-46\pm\sqrt{46^2+23(5355)}}{23}"


Since "x>0," we take


"x=\dfrac{-46+\sqrt{46^2+23(5355)}}{23}"



"x=13"

2.2


"y=1035x^2+4140x-1725, x>0"